# A Lightweight Method for Defending Against UAF Vulnerabilities

### v1

@@ 5. Evaluation
Our experiments mainly evaluate three aspects of LightDE: effectiveness, performance overhead, and scalability in multithreaded applications. First, we experimentally assess LightDE’s effectiveness in defending against real UAF vulnerabilities. Next, we evaluate the additional performance overhead introduced by LightDE using the SPEC CPU2006 benchmark suite (Henning, 2006). We chose the SPEC CPU2006 benchmark suite because existing dangling pointer elimination methods use this suite, making it convenient for us to compare our method with existing methods. Finally, we assess LightDE’s scalability in multithreaded applications using the SPLASH-2X benchmark suite (Group et al., 2011). We believe that heap pointers and global pointers exist for a longer duration and have a wider scope of influence, making them more likely to cause UAF vulnerabilities. Therefore, in our experiments, we not only analyzed the impact on program performance when considering all pointers, including stack, heap, and global pointers, but also separately analyzed the impact on program performance when considering only heap and global pointers. We implemented a prototype system of LightDE based on LLVM-13 (llvm, 2000) and SVF-2.5 (Sui and Xue, 2016; SVF, 2022). All experiments were conducted on a machine running 64-bit Ubuntu 20.04 LTS, equipped with a 13th Gen Intel(R) Core(TM) i7-13700H processor at 2.640GHz and 32GB of RAM.
5.1. Effectiveness
Table 3. Real-world UAF Vulnerabilities Used for Evaluation.
@@ 5.1. Effectiveness
CVE-2024-24260 (CVE-2024-24260, 2024) media-server v1.0.0
CVE-2020-16590 (CVE-2020-16590, 2020) Binutils 2.35
CVE-2018-5747 (CVE-2018-5747, 2018) lrzip 0.631
CVE-2017-12448 (CVE-2017-12448, 2017) Binutils 2.29
To evaluate the effectiveness of LightDE in mitigating UAF vulnerabilities, we applied LightDE to four programs with known UAF vulnerabilities and ran exploitation programs to determine the effectiveness of LightDE in defending against these vulnerabilities. The experimental results show that LightDE was able to prevent the exploitation of UAF vulnerabilities in all four applications, as shown in Table 3.
5.2. Performance on SPEC CPU2006
5.2.1. Statistics
Table 4. Statistics for SPEC CPU2006.
@@ 5.2.1. Statistics
pointers #
400.perlbench 22 253 357 3626 2033 358M 356M 9569K
401.bzip2 6 7 18 3 0 162 144 18
403.gcc 1724 1201 2021962 281490 759291 21M 21M 2938K
429.mcf 4 9 27 10 0 3 3 4
433.milc 61 30 416 37 18 6509 6468 0
444.namd 44 402 21 6 41 1329 1321 56
445.gobmk 51 9 298 65 238 656K 656K 209K
447.dealII 388 1502 28384 12347 49348 24M 24M 14M
450.soplex 269 469 52972 13609 3145 67K 67K 473
453.povray 532 583 284014 45194 302064 2460K 2460K 1747K
456.hmmer 224 330 11754 1 538 1351K 1351K 709K
458.sjeng 10 16 27 7 0 4 0 0
462.libquantum 7 7 0 2 5 95 95 0
464.h264ref 293 582 7446 429 4 176K 176K 828K
470.lbm 2 2 0 4 0 2 2 1
471.omnetpp 2434 2155 831479 204579 49198 312M 266M 451K
473.astar 64 81 138 47 17 4790K 4790K 3049K
482.sphinx3 162 93 5225 100 9 7918K 7918K 4115K
483.xalancbmk 564 5605 253507 192478 443941 126M 126M 476K
@@
Table 4 shows the static data of the SPEC CPU2006 benchmark suite. The second column displays the number of static objects for each project. The seventh column shows the number of objects allocated at runtime. By comparing these two columns, it is evident that a single static object corresponds to multiple runtime objects. The third column shows the number of release points identified by the pointer analysis, while the eighth column shows the number of objects released at runtime. From these two columns, it is also clear that a single release point releases multiple runtime objects. The fourth column shows the number of static heap pointers identified by the pointer analysis, the fifth column shows the number of static global pointers identified by the pointer analysis, and the sixth column shows the number of static stack pointers identified by the pointer analysis. The last column displays the number of dangling pointers eliminated by LightDE. These data show the difference between the program’s static information and its runtime information.
5.2.2. Runtime Overhead
Figure 6. LightDE’s Runtime on SPEC CPU2006.
@@ 5.2.2. Runtime Overhead
Figure 6 shows the runtime overhead introduced by LightDE in each project. When eliminating only heap and global dangling pointers, the geometric mean of the runtime overhead introduced by LightDE is 3.76%. When eliminating all dangling pointers, including heap, stack, and global dangling pointers, the geometric mean of the runtime overhead is 8.66%. It is worth noting that the additional time or memory overhead introduced by LightDE in some projects is far below 0.01%. To calculate the geometric mean, we set the overhead for these projects to 1% in this paper.
@@
As shown in Figure 6, when LightDE eliminates only heap and global dangling pointers, the additional runtime overhead introduced in most projects is very low. However, we also observed significant runtime overhead in the 471.omnetpp and 483.xalancbmk projects. The reason for this is that these projects are memory allocation-intensive applications. These two programs frequently allocate and release memory, requiring frequent maintenance of the hash table and the lock-free queue (Valois, 1995), which leads to significant performance overhead. As shown in Figure 6, when LightDE eliminates stack, heap, and global dangling pointers, the additional runtime overhead increases. This is because our method requires the original thread of the program to store the memory addresses of stack pointers in our designed data structure and notify our thread to remove the corresponding local pointers upon function exit. For example, in the 453.povray and 473.astar projects, the frequent function entry and exit caused relatively high additional runtime overhead. Overall, the additional time overhead introduced by LightDE remains very low.
5.2.3. Memory Overhead
Figure 7. LightDE’s Memory Overhead on SPEC CPU2006.
Figure 8. Runtime Overhead Comparison With Dangling Pointer Elimination Methods.
@@ 5.2.3. Memory Overhead
Figure 7 shows the additional memory overhead introduced by LightDE in each project. The geometric mean of the memory overhead introduced by LightDE is 3.88% when eliminating only heap and global dangling pointers. When eliminating all dangling pointers, including heap, stack, and global dangling pointers, the geometric mean of the memory overhead is 7.26%.
@@
As shown in Figure 7, when considering only heap and global pointers, the additional memory overhead introduced by LightDE is very low in most projects. However, in the projects 447.dealII, 471.omnetpp, and 482.sphinx3, LightDE introduced significantly high memory overhead. We analyzed the reasons for the high memory overhead introduced by LightDE in some projects. The memory overhead of LightDE mainly comes from two sources: the global hash table and the lock-free queue used by the program’s original thread to pass object information to the LightDE thread. LightDE iteratively eliminates dangling pointers in a separate thread. Since the results of pointer analysis can be very inaccurate under certain circumstances, it can lead to LightDE checking too many irrelevant pointers. This ineffective checking can result in excessive time consumption, causing the object information passed from the original thread to the LightDE thread to accumulate in the lock-free queue, which leads to higher memory overhead. When considering stack pointers, we observed that in the 473.astar project, a significantly high memory overhead was introduced. This is due to the frequent function calls in the 473.astar project, which generated a large number of stack pointers, resulting in very high memory consumption. Overall, the memory overhead introduced by LightDE remains very low.
5.2.4. Overhead Comparison
@@
In this section, we compare the runtime and memory overhead of LightDE with all existing methods of eliminating dangling pointers (Lee et al., 2015; Younan, 2015; Van Der Kouwe et al., 2017; Liu et al., 2018; Shen and Dolan-Gavitt, 2020). The additional runtime overhead introduced by existing methods and LightDE is shown in Figure 8. Since FreeSentry only reports results of 7 SPEC CPU2006 projects, we display all the projects reported by FreeSentry in the left image of Figure 8. The ’?’ in the figure indicates that the corresponding overhead for the project was not reported in their method. All the data we used comes from the reports of these methods in their papers.
@@
The geometric mean of the runtime overhead introduced by DangNull in its reported projects is 55%. It is worth noting that DangNull only considers heap pointers and does not take stack pointers and global pointers into account. FreeSentry reported the overhead of eliminating only heap dangling pointers and the overhead of eliminating both heap and stack pointers. Since FreeSentry reported too few SPEC CPU2006 projects, we do not calculate the geometric mean of its runtime overhead. The geometric mean of the runtime overhead introduced by DangSan is 41%. The geometric mean of the runtime overhead introduced by pSweeper is 17.2%. When considering only global and heap pointers, the geometric mean of the runtime overhead introduced by our method is 3.76%. When considering all pointers, the geometric mean is 8.66%. Our method introduces significantly lower runtime overhead compared to existing dangling pointer elimination methods, whether considering stack pointers or not.
@@
When comparing memory overhead, we only compared with the better performing dangling pointer elimination methods DangSan and pSweeper. The geometric mean of the memory overhead introduced by DangSan is 240%, and the geometric mean of the memory overhead introduced by pSweeper is 247.3%. The memory overhead introduced by LightDE is significantly lower than that of existing dangling pointer elimination methods, regardless of whether stack pointers are considered.
Figure 9. LightDE’s Runtime Overhead on SPLASH-2X.
Figure 10. LightDE’s Memory Overhead on SPLASH-2X.
5.3. Scalability on Multi-threaded Applications
@@
We used the SPLASH-2X benchmark suite to evaluate LightDE’s scalability on multithreaded applications. Figures 9 and 10 show the additional memory and runtime overhead introduced by LightDE compared to the original program. We ran all the projects in the SPLASH-2X benchmark suite. We evaluated the overhead of setting each program to run 1, 2, 4, 8, 16, 32, and 64 threads. We evaluated the overhead of using LightDE to eliminate only heap and global dangling pointers, as well as the overhead of using LightDE to eliminate stack, heap, and global dangling pointers.
@@
Figure 9 shows the additional runtime overhead introduced by LightDE in the SPLASH-2X benchmark suite. When eliminating only heap and global dangling pointers, the geometric mean of the runtime overhead caused by LightDE is 1.23%. When eliminating all dangling pointers, including heap, stack, and global dangling pointers, the geometric mean of the runtime overhead is 8.66%. From Figure 9, it can be seen that when only heap and global dangling pointers are eliminated, LightDE introduces almost no additional runtime overhead. In this case, the runtime of the programs protected by LightDE overlaps with the runtime of the unprotected programs. However, when LightDE eliminates stack, heap, and global dangling pointers, it results in significant additional runtime overhead in the ”lu_ncb” and ”radiosity” projects. We analyzed this phenomenon. In our design, the memory addresses of stack pointers are stored in our data structure by the original program threads, which can lead to considerable time overhead. Overall, the additional runtime overhead introduced by LightDE remains very low.
@@
Figure 10 shows the additional memory overhead introduced by LightDE, as well as the memory overhead of the original SPLASH-2X programs. When considering only global and heap pointers, the geometric mean of the memory overhead introduced by LightDE is 2.10%. When considering all pointers, including heap, stack, and global pointers, the geometric mean of the memory overhead is 2.31%. From Figure 10, it can be seen that in most of programs, the memory overhead introduced by LightDE is negligible, regardless of whether stack pointers are considered. However, the projects ”water_nsquared” and ”water_spatial” result in relatively higher memory overhead. In the ”water_nsquared” project, memory overhead increases with the number of threads, indicating that a large number of objects are created by the program’s child threads rather than by the main thread. In the ”water_spatial” project, the memory increase remains almost constant across all thread counts, suggesting that the increase in memory overhead does not correlate with the number of threads. This implies that the memory overhead primarily comes from the main thread, which creates more objects. Overall, the additional memory overhead introduced by LightDE remains very low.
6. Limitations and Future Work

### v4

@@ V Evaluation
Our experiments first evaluate the accuracy improvement of our structure-sensitive pointer analysis compared to Andersen’s pointer analysis. Then, our experiments assess three aspects of LightDE: effectiveness, performance overhead, and scalability on multi-threaded applications. First, we experimentally assess LightDE’s effectiveness in defending against real UAF vulnerabilities. Next, we evaluate the performance overhead introduced by LightDE using the SPEC CPU2006 benchmark suite [32]. Finally, we assess LightDE’s scalability on multi-threaded applications using the SPLASH-2X benchmark suite [33]. We believe that heap pointers and global pointers exist for a longer duration and have a wider scope of influence, making them more likely to cause UAF vulnerabilities. Therefore, in our experiments, we analyzed the impact on program performance when considering all pointers, including stack, heap, and global pointers, as well as the impact when considering only heap and global pointers. In the figure of the experimental results, the label ”with-stack” represents the case where all pointers are considered, while the label ”no-stack” represents the case where only heap and global pointers are considered. We implemented a prototype system of LightDE based on LLVM-13 and SVF [41, 42, 43, 44]. All experiments were conducted on a machine running 64-bit Ubuntu 20.04, equipped with a 13th Gen Intel(R) Core(TM) i7-13700H processor at 2.640GHz and 32GB of RAM.
V-A Accuracy of our Sturcture-sensitivity
Before assessing the performance of LightDE in eliminating dangling pointers, we first evaluate the accuracy improvement of our structure-sensitive pointer analysis method compared to Andersen pointer analysis on the SPEC CPU2006 benchmark suite. Our structure-sensitive pointer analysis method is a two-phase pointer analysis. The first phase of pointer analysis is similar to Andersen’s pointer analysis. The difference between the first stage and the Andersen’s is that our method simultaneously determines the type information of objects while analyzing pointer targets. The second stage of pointer analysis filters out false pointer relationships by using the object type information obtained in the first stage. The main objective of this experiment is to evaluate how many spurious pointing relationships our structure-sensitive pointer analysis reduces compared to Andersen pointer analysis. The first and second columns of Table III show the number of pointer relationships inferred by Andersen’s pointer analysis and the number of pointer relationships inferred by our structure-sensitive pointer analysis for each item, respectively. The third column of the table indicates the number of spurious point-to relationships reduced by our structure-sensitive pointer analysis compared to Andersen pointer analysis for each benchmark. From the data in the last column of the table, we can see that the number of spurious point-to relationships filtered out by our structure-sensitive pointer analysis varies significantly across different benchmarks. For instance, in benchmark 470.lbm, our method shows no accuracy improvement over Andersen pointer analysis, whereas in benchmark 471.omnetpp, our method filtered out more than 50% of the false point-to relationships present in Andersen pointer analysis. The primary reason for this discrepancy lies in the nature of the programs themselves. Our structure-sensitive pointer analysis method filters false point-to relationships based on the types of objects, so our method requires knowledge of the exact offsets used by the instructions to access the objects when filtering false point-to relationships. If the program accesses object fields through variable offsets, our method will be unable to filter out false point-to relationships based on object types. These variable offsets may require complex analysis or be influenced by input data, as illustrated by the example in Figure 5. Therefore, if there are a large number of cases in the program where object fields are accessed via unknown offsets, our method will be unable to filter out false pointing relationships based on object types.
TABLE IV: Real-world UAF Vulnerabilities.
@@ V-A Accuracy of our Sturcture-sensitivity
CVE-2024-24260 [45] media-server v1.0.0
CVE-2020-16590 [46] Binutils 2.35
CVE-2018-5747 [47] lrzip 0.631
CVE-2017-12448 [48] Binutils 2.29
TABLE V: Statistics for SPEC CPU2006.
@@
pointer #
pSweeper
invalidate
pointer #
400.perlbench 22 253 357 3626 2033 358M 356M 18M 421K
401.bzip2 6 7 18 3 0 162 144 18 0
403.gcc 1724 1201 2021962 281490 759291 21M 21M 2938K 186K
429.mcf 4 9 27 10 0 3 3 4 0
433.milc 61 30 416 37 18 6509 6468 0 0
444.namd 44 402 21 6 41 1329 1321 56 0
445.gobmk 51 9 298 65 238 656K 656K 209K 86
447.dealII 388 1502 28384 12347 49348 24M 24M 14M 7293
450.soplex 269 469 52972 13609 3145 67K 67K 473 553
453.povray 532 583 284014 45194 302064 2460K 2460K 1747K 7428
456.hmmer 224 330 11754 1 538 1351K 1351K 709K 0
458.sjeng 10 16 27 7 0 4 0 0 0
462.libquantum 7 7 0 2 5 95 95 0 0
464.h264ref 293 582 7446 429 4 176K 176K 828K 961
470.lbm 2 2 0 4 0 2 2 1 0
471.omnetpp 2434 2155 831479 204579 49198 312M 266M 4694k 13K
473.astar 64 81 138 47 17 4790K 4790K 3049K 104
482.sphinx3 162 93 5225 100 9 7918K 7918K 4115K 762
483.xalancbmk 564 5605 253507 192478 443941 126M 126M 18M 79K
V-B Effectiveness
To evaluate the effectiveness of LightDE in mitigating UAF vulnerabilities, we applied LightDE to four programs with known UAF vulnerabilities. We then ran exploitation programs to determine LightDE’s effectiveness in defending against these vulnerabilities. The experimental results show that LightDE was able to prevent the exploitation of UAF vulnerabilities in all four applications, as shown in Table IV.
V-C Performance on SPEC CPU2006
V-C1 Statistics
@@
Table V shows the static data of the SPEC CPU2006 benchmark suite. The second column displays the number of static objects for each project. The seventh column shows the number of objects allocated at runtime. By comparing these two columns, it is evident that a single static object corresponds to multiple runtime objects. The third column shows the number of release points identified by the pointer analysis, while the eighth column shows the number of objects released at runtime. From these two columns, it is also clear that a single release point releases multiple runtime objects. The fourth column shows the number of static heap pointers identified by the pointer analysis, the fifth column shows the number of static global pointers identified by the pointer analysis, and the sixth column shows the number of static stack pointers identified by the pointer analysis. The ninth column displays the number of dangling pointers eliminated by LightDE. The last column of the table shows the number of dangling pointers eliminated by pSweeper [24] in each project. Based on the last two columns of the table, it can be observed that LightDE eliminates significantly more dangling pointers than pSweeper.
Figure 8: LightDE’s Runtime Overhead on SPEC CPU2006.
Figure 9: Runtime Overhead Comparison. ’?’ indicates the corresponding overhead was not reported in the method.
Figure 10: LightDE’s Memory Overhead on SPEC CPU2006.
V-C2 Runtime Overhead
@@
Figure 8 shows the runtime overhead introduced by LightDE in each project. When only the heap and global pointers are considered, LightDE introduces an additional geometric mean runtime overhead of 3.76% compared to the unprotected original program. When considering all pointers, including heap, stack, and global pointers, LightDE introduces an additional geometric mean runtime overhead of 8.66% compared to the unprotected original program. It is worth noting that the time or memory overhead introduced by LightDE in some projects is far below 0.01%. To calculate the geometric mean, we set the overhead for these projects to 1% in this paper.
@@
As shown in Figure 8, when LightDE considers only heap and global pointers, the runtime overhead introduced in most projects is very low. However, we also observed significant runtime overhead in the 471.omnetpp and 483.xalancbmk projects. The reason for this is that these projects are memory allocation-intensive applications. These two projects frequently allocate and release memory, requiring frequent maintenance of the hash table and the lock-free queue, which leads to significant performance overhead. As shown in Figure 8, when LightDE considers all pointers, including stack, heap, and global pointers, the runtime overhead increases. This is because our method requires the original thread of the program to store the memory addresses of stack pointers in our designed data structure and notify our thread to remove the corresponding local pointers upon function exit. For example, in the 453.povray and 473.astar projects, the frequent function entry and exit caused relatively high runtime overhead.
V-C3 Memory Overhead
@@
Figure 10 shows the memory overhead introduced by LightDE in each project. When considering only heap and global pointers, LightDE introduces a geometric mean memory overhead of 3.88% compared to the unprotected original program. When considering all pointers, including heap, stack, and global dangling pointers, LightDE introduces a geometric mean memory overhead of 7.26%.
@@
As shown in Figure 10, when considering only heap and global pointers, the memory overhead introduced by LightDE is very low in most projects. However, in the projects 447.dealII, 471.omnetpp, and 482.sphinx3, LightDE introduced significantly high memory overhead. We analyzed the reasons for the high memory overhead introduced by LightDE in these projects. The memory overhead of LightDE mainly comes from the inability to timely release objects that are awaiting release. Since the results of pointer analysis can be very inaccurate under certain circumstances, it can lead to LightDE thread checking too many irrelevant pointers. This ineffective checking leads to excessive time consumption, causing the objects awaiting release to accumulate in the shared queue between the original thread and the LightDE thread, resulting in higher memory overhead. When considering stack pointers, we observed that in the 473.astar project, a significantly high memory overhead was introduced. This is due to the frequent function calls in the 473.astar project, which generated a large number of stack pointers, resulting in very high memory consumption.
Figure 11: LightDE’s Runtime Overhead on SPLASH-2X.
Figure 12: LightDE’s Memory Overhead on SPLASH-2X.
V-C4 Overhead Comparison
@@
In this section, we compare the runtime and memory overhead of LightDE with existing methods of eliminating dangling pointers [22, 7, 23, 24, 25]. The runtime overhead introduced by existing methods and LightDE is shown in Figure 9. Since FreeSentry only reports results of 7 SPEC CPU2006 projects, we display all the projects reported by FreeSentry in the left image of Figure 9.
@@
The geometric mean of the runtime overhead introduced by DangNull in its reported projects is 55%. It is worth noting that DangNull only considers heap pointers and does not take stack pointers and global pointers into account. FreeSentry reported the overhead of eliminating only heap dangling pointers and the overhead of eliminating both heap and stack pointers. Since FreeSentry reported too few SPEC CPU2006 projects, we do not calculate the geometric mean of its runtime overhead. The geometric mean of the runtime overhead introduced by DangSan is 41%. The geometric mean of the runtime overhead introduced by pSweeper is 17.2%. When considering only global and heap pointers, the geometric mean of the runtime overhead introduced by our method is 3.76%. When considering all pointers, the geometric mean of the runtime overhead is 8.66%. Our method introduces significantly lower runtime overhead compared to existing dangling pointer elimination methods, whether considering stack pointers or not.
@@
We only compared memory overhead with the better performing dangling pointer elimination methods DangSan and pSweeper. The geometric mean of the memory overhead introduced by DangSan is 240%, and the geometric mean of the memory overhead introduced by pSweeper is 247.3%. The memory overhead introduced by LightDE is significantly lower than that of existing dangling pointer elimination methods, regardless of whether stack pointers are considered.
During our experiments, we had a question: Is LightDE’s low runtime overhead mainly due to the fact that it determines and stores pointer relationships during compilation, or is it due to multithreading? By comparing the specific processes of eliminating dangling pointers in pSweeper and our method, LightDE, we can reaches the conclusion that the reduced runtime overhead of LightDE is primarily due to determining and storing pointer relationships at compile time. pSweeper requires the original thread to record the information of allocated objects and the memory addresses of pointers during program execution, and the original thread records the released object information into a specific data structure when an object is released. And pSweeper uses a separate thread to scan the pointers recorded in the global data structure to check whether they point to released objects. LightDE similarly requires the original thread to record the information of allocated objects during program execution, and when an object is released, the original thread stores the information of the released object into a specific data structure. LightDE similarly uses a separate thread to check whether the pointers identified by static pointer analysis still point to the released object. Since the separate threads that pSweeper and LightDE run on idle CPUs do not affect the program’s runtime overhead, the analysis of runtime overhead mainly compares the operations of pSweeper and LightDE on the original threads. The only difference between pSweeper and LightDE on the original thread is that pSweeper requires the original thread to record the memory addresses of pointers during program execution. Therefore, we can conclude that LightDE’s reduced runtime overhead is mainly due to determining and storing the pointer relationships at compile time.
V-D Scalability on Multi-threaded Applications
@@
We used the SPLASH-2X benchmark suite to evaluate LightDE’s scalability on multi-threaded applications. Figures 11 and 12 show the memory and runtime overhead introduced by LightDE compared to the original program. We ran all the projects in the SPLASH-2X benchmark suite. We evaluated the overhead of setting each program to run 1, 2, 4, 8, 16, 32, and 64 threads. We evaluated the overhead of using LightDE to eliminate only heap and global dangling pointers, as well as the overhead of using LightDE to eliminate all dangling pointers, including heap, stack, and global dangling pointers.
@@
Figure 11 shows the runtime overhead introduced by LightDE on the SPLASH-2X benchmark suite. When considering only heap and global pointers, the geometric mean of the runtime overhead caused by LightDE is 1.23%. When considering all pointers, including heap, stack, and global dangling pointers, the geometric mean of the runtime overhead is 8.66%. From Figure 11, it can be seen that when only heap and global pointers are considered, LightDE introduces almost no runtime overhead. In this case, the runtime of the programs protected by LightDE overlaps with the runtime of the unprotected programs. However, when LightDE considers stack, heap, and global dangling pointers, it results in significant runtime overhead in the ”lu_ncb” and ”radiosity” projects. We analyzed this phenomenon. In our design, the memory addresses of stack pointers are stored in our data structure by the original program threads. This indicates that these two projects experience frequent function entries and exits, causing the original program threads to frequently store the memory addresses of stack pointers.
@@
Figure 12 shows the memory overhead introduced by LightDE on the SPLASH-2X benchmark suite. When considering only global and heap pointers, the geometric mean of the memory overhead introduced by LightDE is 2.10%. When considering all pointers, including heap, stack, and global pointers, the geometric mean of the memory overhead is 2.31%. From Figure 12, it can be seen that in most projects, the memory overhead introduced by LightDE is negligible, regardless of whether stack pointers are considered. However, the projects ”water_nsquared” and ”water_spatial” result in relatively higher memory overhead. In the ”water_nsquared” project, memory overhead increases with the number of threads, indicating that a large number of objects are created by the program’s child threads rather than by the main thread. In the ”water_spatial” project, the memory increase remains almost constant across all thread counts, suggesting that the increase in memory overhead does not correlate with the number of threads. This implies that the memory overhead primarily comes from the main thread, which creates more objects.
VI Limitations and Future Work